Gucchi:

A coin is flipped 150 times. The results of the experiment are shown in the following table: Heads Tails 63 87 Which of the following statements best describes the experimental probability of getting heads? It is equal to the theoretical probability. It is 8% lower than the theoretical probability. It is 8% higher than the theoretical probability. The experimental probability cannot be concluded from the data in the table.

Gucchi:

1 attachment
Gucchi:

Am I right with B?

Gucchi:

@imagine

Gucchi:

@hhanan

Gucchi:

@astrid1

Gucchi:

@tetsxpreme

Florisalreadytaken:

A coin is flipped `150 times`. the theoretical is \(50/50\) as the coin has got 2 sides its either heads or tails the experimental probability is Heads Tails 63 87 were asked to find the probability of pulling HEADS -- thus, we can write: \[ \frac{63}{150} \overset{/3}\Rightarrow \frac{21}{50} \] to turn that into a percentage, we have to multiply by 2: \[ \frac{21}{50} \times 2 \Rightarrow \frac{42}{100} \] that means that the probability we are pulling heads is \( 42\% \) that is indeed \(8\% \) lower than the theoretical probability. Well done!

Gucchi:

THANK YOU, wasnt sure anyone was going to answer